OpenStudy (anonymous):

I will fan and medal! simplify: ((a^4)/(2a^2-4a^4))*((8a^3+1)/(2a+4a^2))

OpenStudy (anonymous):

ugh i wish you had typed it out properly :(

OpenStudy (anonymous):

you have a calculator?

jimthompson5910 (jim_thompson5910):

@Jdosio \[\Large \frac{a^4}{2a^2-4a^4} \times \frac{8a^3+1}{2a+4a^2}\]

OpenStudy (anonymous):

or go to www.mathway.com then type in this equation and hit solve

OpenStudy (welshfella):

try factoring the individual expressions first

OpenStudy (anonymous):

if you have any choices that would help too.

jimthompson5910 (jim_thompson5910):

hint: factor \[\Large \frac{a^4}{2a^2-4a^4} \times \frac{8a^3+1}{2a+4a^2}\] \[\Large \frac{a^4}{2a^2(1-2a^2)} \times \frac{(2a+1)(4a^2-2a+1)}{2a(1+2a)}\]

jimthompson5910 (jim_thompson5910):

I used the sum of cubes formula to factor \(\Large 8a^3 + 1\)

OpenStudy (welshfella):

you can now cancel out the 2a + 1 on the right most fraction and the left one simplifies to a^2 ------ 2(1 - 2a^2)